the magnetic induction field due to a magnetic north pole of strength 10am and 5cm away from the pole in weber/msquare

Grade:

1 Answers

Khimraj
4 years ago
 
Magnitude of magnetic induction due to a pole of strength m at a distance r from it is
B = (\mu0 / 4\pi)(m / r2) ,     \mu0 / 4\pi = 10–7 H.
m = 10 Am,  r = 5 cm = 5 x 10–2 m
So, magnitude of magnetic induction, B = (10–7)(10 / 25 x 10–4)  =  4 x 10–12 Wb / m2.